The threshold wavelength of a photosensitive metal is 275nm. Calculate the maximum velocity of the photoelectrons ejected when light of 180nm incident on it

Grade:12

1 Answers

Nikhil
5 years ago
hf-hf(0)=K.Emax
1240/180-1240/275=K.Emax in eV
2.3eV=K.Ein eV
2.3×1.6×10-19=3.68×10-19V =potential and thw max speed of electron